OpenStudy (anonymous):

Given the number, two-fifths, generate its equivalent forms as a fraction, a decimal, and a percent. In addition, give a real world example in which each form might be used.

OpenStudy (anonymous):

@shinebrightlikeadimon ???

OpenStudy (anonymous):

@Michele_Laino he os she is really smart im sorry but idk this one

OpenStudy (michele_laino):

we have: \[\frac{ 2 }{ 5 }=\frac{ 4 }{ 10 }=0.4\] furthermore: \[\frac{ 2 }{ 5 }=\frac{ 4 }{ 10 }=\frac{ 40 }{ 100}= 40 percent\]

OpenStudy (michele_laino):

for example 0.4 can be the diameter of a stylus pen, whereas 40% can be a percentage discount to the supermarket

OpenStudy (anonymous):

so is 4/10 the equivalent fraction ?

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

okay thank you :)

OpenStudy (michele_laino):

also 2/5 is an equivalent fraction

OpenStudy (michele_laino):

thank you!

OpenStudy (anonymous):

yw